# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v3

@@ 4. CataractBot System Design
Here, we describe the CataractBot system (Figure 3), including its key components–input language and modality, response generation and verification, and escalation and reminders–and provide a detailed account of its implementation.
@@ 4.1.1. Input Language and Modality
Sankara Eye Hospital in Bangalore (in the state of Karnataka), referred to as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. Analysis of a 2011 census highlights Bangalore as one of India’s most linguistically diverse cities (Office of the Registrar General & Census Commissioner, 2011). To accommodate this diverse patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). As the expert may not be proficient in all five languages, their interactions with the bot are exclusively in English. Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language. Among these messages is an option allowing users to modify their preferred language in future. To accommodate different levels of literacy, including literate, semi-literate, and illiterate users, CataractBot supports both text and speech inputs. For every voice message, CataractBot sends both a text response and a spoken version of the text as an audio message.
Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot provides a set of three frequently asked questions along with the welcome message. Additionally, every bot’s response includes a “What to do next?” prompt, offering three related questions based on the preceding query (Figure 2F). This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves the user time.
@@ 4.1.2. Response Generation, Verification and Icons
Upon receiving a message, CataractBot classifies it as a medical question (e.g., dos and donts before and after surgery), logistical question (e.g., scheduling or insurance related), small talk, or ‘other’, and provides a response in real-time.
For medical and logistical questions, the bot strictly employs the KB curated by the hospital’s medical team to generate an appropriate response.
@@
In instances where the custom KB lacks an answer, the bot responds with a template “I don’t know” response.
@@
For medical questions, the operating doctor of the respective patient receives a message comprising of the question asked, the bot’s response, and patient demographics (Figure 1A).
@@
The CataractBot system automatically combines the bot’s initial answer with the doctor-provided correction to frame a new response, which is delivered to the user with a green tick icon
@@
Similar to the doctor’s workflow, the patient coordinators verify and provide corrections to the bot’s response for logistical questions.
We deliberated between displaying unverified responses in real-time and presenting only verified responses post-expert verification and correction.
Through our
formative study and discussions with hospital staff, it became clear that providing real-time responses was crucial.
@@ 4.1.3. Escalation Mechanism
4.1.4. KB Update Process
To minimize experts’ labor, we use the expert-provided edits to update the KB, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future. However, certain responses, such as those specific to individual patients (e.g., “When to reach the hospital for my surgery?”) must not be added to the KB as they are not generalizable. We appointed a senior cataract surgeon as the ‘KB update expert’, who received a spreadsheet via email at 8pm daily. Their task involves reviewing each row and determining whether the information should be added to the KB (by responding with a ‘Yes’/‘No’ in the ‘Should Update KB?’ column), and if so, modifying as needed the ‘Final Answer for KB’ column, containing the bot’s updated answer based on the expert’s correction. At 3 am daily, the system extracts ‘Question’ and ‘Final Answer for KB’ data from all rows marked ‘Yes’, and append these to the KB in the ‘expert-FAQ’ document. Additionally, we prompt the LLM to prioritize this ‘expert-FAQ’ document within the KB. This ensures that CataractBot improves with these updates, gradually enhancing its accuracy over time.
@@ 4.2. Implementation Details
The CataractBot system relies on these five components (Figure 3): (a) LLM: for response generation, (b) Vector Database: for storing and retrieving the custom KB, (c) Language Technologies: for translation and transcription, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
@@
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). The custom KB comprises of various cataract surgery and hospital-specific documents, including Consent Form, SOP (Standard Operating Procedure), FAQs, Pre- and Post-Operative Guidelines, and Hospital Information. These documents are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question, GPT-4 classifies the query type. For a medical/logistical question, CataractBot employs a RAG (Lewis et al., 2020) approach. This involves performing a vector search on the KB to extract the three most relevant data chunks related to the query. GPT-4 is then prompted (full prompt available in Appendix A) to generate an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message.
@@
The prompt underwent several iterations. To validate its effectiveness, a doctor and a patient coordinatorrecorded 153 questions posed by patients/attendants scheduled for cataract surgery over a week. We used our GPT-4 prompt to obtain answers to these questions from the custom KB, and subsequently the generated responses were evaluated by the same doctor and coordinator. This not only helped refine the LLM prompt, but also led the medical staff to improve existing documents and contribute additional documents to enrich the KB. This process was iterated thrice.
@@
Additionally, we utilize GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 executed this task well. Regarding language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, our system adhered to the standard approach (ait, [n. d.]) of translating Indic languages into English for input using Azure AI Translator. It then translated GPT-4’s English responses back into the respective Indic languages for output. To facilitate speech as input and audio as output, CataractBot uses Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.) Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and KB update expert) get logged in a cloud storage for further analysis.
@@ 5. Study Design
To evaluate the effectiveness of this LLM-powered experts-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 at our collaborators’ institute, Sankara Eye Hospital. This is one of the leading tertiary eye care and teaching institutions in Bangalore, India. It typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital. None of the study participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation, in accordance with hospital norms.
@@ 5.1. Procedure for Patients and Attendants
All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages. (Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while seven days post-surgery interviews were conducted either in person at the hospital or through pre-scheduled telephone calls. This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with the procedure itself taking 45 minutes. Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time. Each interview typically spanned 20 to 40 minutes and was audio-recorded with the participant’s consent.
@@ 5.2. Procedure for Doctors and Coordinators
In addition to patients and attendants, we recruited four doctors and two patient coordinators. Among the doctors, three served as operating doctors, and one as both an escalation doctor and KB update expert. Both patient coordinators specialize in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator. The escalation doctor and escalation coordinator held senior positions in the hospital and carried administrative responsibilities. To ensure preparedness, all participating experts attended one of two one-hour training sessions conducted by three researchers. During these sessions, CataractBot was demonstrated, and they were onboarded onto the platform. The researchers assumed patient roles, asking medical and logistical questions for the doctors and coordinators to verify and edit. The first session revealed a few bugs, and feedback was collected to enhance the bot. For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in ‘tagged’ WhatsApp messages (Figure 1E). After using CataractBot for two weeks or longer, the participating experts were interviewed by the first author. These interviews followed a semi-structured format, focusing on the bot’s usability, its integration into their workflow, and its impact on interactions with patients. The interviews were audio-recorded, conducted in English, in-person, and typically lasted 45-60 minutes.

### v4

@@ 4. CataractBot System Design
Here, we describe the key components of the CataractBot system (Figure 3)–input language and modality, response generation and verification, escalation and reminders–and provide a detailed account of its implementation.
@@ 4.1.1. Input Language and Modality
Sankara Eye Hospital in Bangalore (in the state of Karnataka), referred to as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. Analysis of a 2011 census highlights Bangalore as one of India’s most linguistically diverse cities (Office of the Registrar General & Census Commissioner, 2011). To accommodate this varying patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). As the expert may not be proficient in all five languages, their interactions with the bot are exclusively in English. Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language. Among these messages is an option allowing users to modify their preferred language in future. To engage with users with different levels of literacy, including literate, semi-literate, and illiterate users, CataractBot supports both text and speech inputs. For every voice message, CataractBot sends both a text response and a spoken version of the text as an audio message.
Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot provides a set of three frequently asked questions along with the welcome message. Additionally, every response includes a “What to do next?” prompt, offering three related questions based on the preceding query (Figure 2F). Users can select one of the suggested questions or pose a different query to continue the conversation. This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves time.
@@ 4.1.2. Response Generation, Verification and Icons
Upon receiving a message, CataractBot classifies it as a medical question (e.g., dos and donts before and after surgery), logistical question (e.g., scheduling or insurance related), small talk (e.g., greetings or expressions of gratitude), or ‘other’, and provides a response in real-time.
For medical and logistical questions, the bot strictly employs the knowledge base curated by the hospital’s medical team to generate an appropriate response.
@@
In instances where the custom knowledge base lacks an answer, the bot responds with a template “I don’t know” response.
@@
For medical questions, the patient’s operating doctor receives a message comprising of the question asked, the bot’s response, and patient demographics (Figure 1A).
@@
The system automatically combines the bot’s initial answer with the doctor-provided correction to frame a new response, which is delivered to the user with a green tick icon
@@
Similar to the doctors’ workflow, patient coordinators verify and provide corrections to the bot’s responses for logistical questions, and have the option to ‘Send to Doctor’ if a medical question is misclassified.
We deliberated between displaying unverified responses in real-time or only verified responses post-expert verification and correction.
Through our formative study and discussions with hospital staff, it became clear that providing real-time responses was crucial.
@@ 4.1.3. Escalation Mechanism
4.1.4. Knowledge Base Update Process
To minimize experts’ labor, we use the expert-provided edits to update the knowledge base, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future. However, certain responses, such as those specific to individual patients (e.g., “When to reach the hospital for my surgery?”) must not be added to the knowledge base as they are not generalizable. We appointed a senior cataract surgeon as the ‘knowledge base expert’, who received a spreadsheet via email at 8pm daily. Their task involves reviewing each row, which consists of question-answer pairs, and determining whether the information should be added to the knowledge base (by responding with a ‘Yes’/‘No’ in the ‘Should Update Knowledge Base?’ column), and if so, modifying as needed the ‘Final Answer for Knowledge Base’ column, containing the bot’s updated answer based on the expert’s correction. At 3 am daily, the system extracts ‘Question’ and ‘Final Answer for Knowledge Base’ data from all rows marked ‘Yes’, and append these to the knowledge base in the ‘expert-FAQ’ document. Additionally, we prompt the LLM to prioritize this ‘expert-FAQ’ document within the knowledge base. This ensures that CataractBot improves with these updates, gradually enhancing its accuracy over time.
@@ 4.2. Implementation Details
The CataractBot system relies on these five components (Figure 3): (a) LLM: for response generation, (b) Vector Database: for storing and retrieving the custom knowledge base, (c) Language Technologies: for translation and transcription, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
@@
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). The custom knowledge base comprises of various cataract surgery and hospital-specific documents, including Consent Form, SOP (Standard Operating Procedure), FAQs, Pre- and Post-Operative Guidelines, and Hospital Information. These documents are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question, GPT-4 classifies the query type. For a medical/logistical question, CataractBot employs a retrieval-augmented generation (Lewis et al., 2020) approach. This involves performing a vector search on the knowledge base to extract the three most relevant data chunks related to the query. GPT-4 is then prompted (full prompt available in Appendix A) to generate an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message.
@@
The prompt underwent several iterations. To validate its effectiveness, a doctor and a patient coordinatorrecorded 153 questions posed by patients/attendants scheduled for cataract surgery over a week. We used our GPT-4 prompt to obtain answers to these questions from the custom knowledge base, and subsequently the generated responses were evaluated by the same doctor and coordinator. This not only helped refine the LLM prompt, but also led the medical staff to improve existing documents and contribute additional documents to enrich the knowledge base. This process was iterated thrice.
@@
Additionally, we utilize GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 executed this task well. Regarding language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, our system adhered to the standard approach (ait, [n. d.]) of translating Indic languages (Hindi, Kannada, Tamil, and Telugu) into English for input using Azure AI Translator. It then translated GPT-4’s English responses back into the respective Indic languages for output. To facilitate speech as input and audio as output, CataractBot uses Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.) Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and knowledge base expert) get logged in a cloud storage for further analysis.
@@ 5. Study Design
To evaluate the effectiveness of this LLM-powered experts-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 at our collaborators’ institute, Sankara Eye Hospital. This is one of the leading tertiary eye care and teaching institutions in Bangalore, India. It typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital.. None of the study participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation, in accordance with hospital norms.
@@ 5.1. Procedure for Patients and Attendants
All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages. (Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while post-surgery interviews were conducted either in person at the hospital or through pre-scheduled telephone calls. This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with the procedure itself taking 45 minutes. Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time. Each interview typically spanned 20 to 40 minutes and was audio-recorded with the participant’s consent.
@@ 5.2. Procedure for Doctors and Coordinators
We recruited four doctors and two patient coordinators. Three doctors served as operating doctors, and one as both an escalation doctor and knowledge base expert. Both patient coordinators specialize in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator. The escalation doctor and escalation coordinator held senior positions in the hospital and carried administrative responsibilities. All participating experts attended one of two one-hour training sessions conducted by three researchers. During these sessions, CataractBot was demonstrated, and experts were onboarded. The researchers assumed patient roles, asking medical and logistical questions for the doctors and coordinators to verify and edit. The first session revealed a few bugs, and feedback was collected to enhance the bot. For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in ‘tagged’ WhatsApp messages (Figure 1E). After using CataractBot for two weeks or longer, the participating experts were interviewed by the first author. These interviews followed a semi-structured format, focusing on the bot’s usability, its integration into their workflow, and its impact on interactions with patients. The interviews were audio-recorded, conducted in English, in-person, and typically lasted 45-60 minutes.